OpenStudy (anonymous):

How many grams of NaI would be used to produce a 2.0 M solution with a volume of 10.00 L?

OpenStudy (aaronq):

the molarity formula: \(Molarity=\dfrac{n}{L}=\dfrac{\dfrac{m}{MW}}{L}=\dfrac{m}{L*MW}\) MW=molar mass m=mass L=liters of solution n=moles

OpenStudy (aaronq):

i'll even rearrange it for ya, m=M*MW*L M= molarity